# Indirect forces in disc-planet interaction

### v1

@@ C.1 Lindblad vs corotation torque separation
For $p=0$ one can see both the oscillatory part of $\mathrm{d}T_{\mathrm{p,id}}/\mathrm{d}R$ roughly outside of the $m=1$ OLR and the peak at the corotation ($R=R_{\mathrm{p}}$). Note that the corotation peak is not particularly narrow, and to obtain $T_{\mathrm{c,id}}$ we integrate $\mathrm{d}T_{\mathrm{p,id}}/\mathrm{d}R$ up to the somewhat arbitrarily selected radius $R=1.2R_{\mathrm{p}}$, roughly where $\mathrm{d}T_{\mathrm{p,id}}/\mathrm{d}R$ changes sign. The remaining integral of $\mathrm{d}T_{\mathrm{p,id}}/\mathrm{d}R$, from $R=1.2R_{\mathrm{p}}$ and above, is then identified with $T_{\mathrm{L,id}}$. This procedure is different from the one employed in Fairbairn2025a to split the direct planetary torque on the disc into the corotation and Lindblad contributions — measuring the jump of the disc AMF $F_{J}$ at the corotation resonance. This is because here we are interested in the indirect effect of the disc on the planet, and not vice versa.
@@
Note that for $p=0$ the peak value of $\mathrm{d}T_{\mathrm{p,id}}/\mathrm{d}R$ at corotation increases as $h_{\mathrm{p}}$ decreases, which is caused by the decreasing softening length (linearly scaling with $h_{\mathrm{p}}$) of the planetary potential that contributes to the density perturbation at the corotation. For that reason the strength of the indirect corotation contribution $T_{\mathrm{c,id}}$ should depend on $h_{\mathrm{p}}$, with the scaling $T_{\mathrm{c,id}}\propto h_{\mathrm{p}}^{-1}$ (GT79), which is indeed the case (see Section 5.1).
For $p=1.5$ (Figure 11b) the peak at the corotation is absent, in agreement with the theoretical expectation (GT79; GT80) for the lack of corotation torque in a $p=1.5$ disc, which has no vortensity gradient. Thus, in this case we assign the full indirect torque due to the disc to $T_{\mathrm{L,id}}$.
Figure 11 also clearly shows that in the framework of our linear calculation the inner disc provides essentially no contribution to $T_{\mathrm{p,id}}$, consistent with analytical expectations. Indeed, $T_{\mathrm{p,id}}$ is determined by the $m=1$ Fourier harmonic $\delta\Sigma_{1}=\Sigma_{1}$ of the surface density perturbation (see equation (36)), which is extremely small in the inner disc for $M_{\mathrm{p}}\ll M_{\mathrm{th}}$ according to Figure 2b. We note in this regard that Crida2025 did find a considerable non-zero (albeit rapidly oscillating) contribution to $T_{\mathrm{p,id}}$ from the inner disc. This may be somehow related to the higher $M_{\mathrm{p}}=0.08M_{\mathrm{th}}$ that they considered; our $M_{\mathrm{p}}=0.3M_{\mathrm{th}}$ calculation clearly shows small but non-zero $\Sigma_{1}$ in the inner disc (see Figure 3b), likely caused by the growth of disc eccentricity in the inner disc, see Section 3.2.
Both panels of Figure 11 make it clear that excitation of the Lindblad torque $T_{\mathrm{L,id}}$ is not concentrated in the immediate vicinity of $R_{\mathrm{L}}$, as expected from the classical analysis of (GT79; GT80) and mentioned in Section 5.1. In reality, near and outside the $m=1$ OLR, $\mathrm{d}T_{\mathrm{p,id}}/\mathrm{d}R$ exhibits slowly decaying oscillations and $T_{\mathrm{L,id}}$ converges to its final value rather far [^10] The non-locality of tidal coupling at Lindblad resonances is also the reason behind the negative torque density phenomenon studied in Dong2011 and RP12. from the planet. The origin of these oscillations is analogous to that of the binary torque oscillations in circumbinary discs (Cimerman2024a) and can again be traced to the global structure of the $m=1$ perturbation in the disc (see Figure 1).
@@ C.2 Direct $m=1$ Lindblad torque $T_{\mathrm{L,d}}$
To compute the direct $m=1$ Lindblad torque component $T_{\mathrm{L,d}}$ we will use the fact that the direct planetary torque on the disc is equal to the direct disc torque on the planet (e.g. see equation (30)). This allows us to use the torque calculation machinery developed by GT79, see equations (A7)-(A10) in the Appendix of that work. However, we need to generalize their results, since GT79 computed the torque exerted by the potential perturbation component $\varphi_{1}$ on the surface density perturbation $\sigma_{1}$ generated by the same potential perturbation $\varphi_{1}$; note that here the subscripts in $\sigma_{1}$ and $\varphi_{1}$ refer to first-order perturbations, not $m=1$ Fourier components. In our case, $\sigma_{1}$ is due to the full perturbing potential — i.e. the sum of the direct and indirect planetary potentials – but the calculation of $T_{\mathrm{L,d}}$ must account only for the torque due to the direct planetary potential acting on this $\sigma_{1}$.
Let us consider a general case when the $\Sigma$ perturbation is caused by a perturbing potential (its $m$-th azimuthal harmonic) $\Phi_{\Sigma}$, while the force enacting the torque is the gradient of some other potential $\Phi_{\mathrm{T}}$. This means that in the equation (A7) of GT79 we associate $\varphi_{1}$ with $\Phi_{\mathrm{T}}$, while $\sigma_{1}$ is induced by $\Phi_{\Sigma}$. Propagating this through the derivation in the equations (A7)-(A10) of GT79, we find the corresponding Lindblad torque at $m$-th OLR to be
@@
In our case the perturbation of $\Sigma$ is caused by the full planetary potential, including its indirect part, $\Phi_{\Sigma}=-(GM_{\mathrm{p}}/R_{\mathrm{p}})[b^{(1)}_{1/2}(\alpha)-\alpha]$, $\alpha=R/R_{\mathrm{p}}$ for $m=1$ (GT80). To find the direct part $T_{\mathrm{L,d}}$ of the $m=1$ Lindblad torque, we take $\Phi_{\mathrm{T}}$ to be given by only the direct part of this expression, $\Phi_{\mathrm{T}}=-(GM_{\mathrm{p}}/R_{\mathrm{p}})b^{(1)}_{1/2}(\alpha)$. Plugging this into (48) and setting $m=1$, $p=1.5$ (to express $\Sigma(R_{1})$ through $\Sigma_{\mathrm{p}}$ and enable comparison in Figure 12) and $R_{\mathrm{L}}=R_{1}=2^{2/3}R_{\mathrm{p}}$, we obtain
@@
which is independent of $h_{\mathrm{p}}$, in lieu of the definition (15). This is in agreement with the classical theory (GT79; GT80) predicting that the torque excited at an individual Lindblad resonance should be independent of $c_{\mathrm{s}}$ and $h_{\mathrm{p}}$.
@@ C.3 Indirect Lindblad torque $T_{\mathrm{L,id}}$
To find $T_{\mathrm{L,id}}$ we cannot use the symmetry property of the direct torques that we exploited for computing $T_{\mathrm{L,d}}$, see the discussion around equations (31) and (32). Instead, we note that our equation (36) looks identical to the initial torque expression (A7) of GT79 if we set $\varphi_{1}=\Phi_{\mathrm{T}}=(GM_{\mathrm{p}}R_{\mathrm{p}}/R^{2})$; this implies that the subsequent result (48) can still be used to compute $T_{\mathrm{L,id}}$ if we adopt this ‘effective’ form of $\Phi_{\mathrm{T}}$ (Fairbairn2025a), keeping $\Phi_{\Sigma}$ as above. Performing this procedure for a $p=1.5$ disc, we arrive at
@@ C.4 Lindblad torque estimates: verification
We verify our analytical calculation of $T_{\mathrm{L,d}}$ and $T_{\mathrm{L,id}}$ in Figure 12 using our numerical linear calculations for $p=1.5$ (when $T_{\mathrm{c,id}}=0$) and $q=0$, finding good agreement with the analytical results given by equations (50) and (51). The variation of these torque components with $h_{\mathrm{p}}$ is rather minor, as expected. The slight differences for lower $h_{\mathrm{p}}$ are likely caused by the non-locality of disc-planet coupling at the $m=1$ OLR, as noted earlier. Our present results, as well as those of KP93, RP12 and Fairbairn2025a, suggest that the simplifying assumption of the locality of Lindblad torque excitation near the OLR, which was adopted in GT79; GT80, may result in some minor inaccuracies.

### v3

@@ C.1 Lindblad vs corotation torque separation
For $p=0$ one can see both the oscillatory part of $\mathrm{d}T_{\mathrm{p,id}}/\mathrm{d}R$ roughly outside of the $m=1$ OLR and the peak at the corotation ($R=R_{\mathrm{p}}$). Note that the corotation peak is not particularly narrow, and to obtain $T_{\mathrm{c,id}}$ we integrate $\mathrm{d}T_{\mathrm{p,id}}/\mathrm{d}R$ up to the somewhat arbitrarily selected radius $R=1.2R_{\mathrm{p}}$, roughly where $\mathrm{d}T_{\mathrm{p,id}}/\mathrm{d}R$ changes sign. The remaining integral of $\mathrm{d}T_{\mathrm{p,id}}/\mathrm{d}R$, from $R=1.2R_{\mathrm{p}}$ and above, is then identified with $T_{\mathrm{L,id}}$. This procedure is different from the one employed in Fairbairn and Rafikov (2025) to split the direct planetary torque on the disc into the corotation and Lindblad contributions — measuring the jump of the disc AMF $F_{J}$ at the corotation resonance. This is because here we are interested in the indirect effect of the disc on the planet, and not vice versa.
@@
Note that for $p=0$ the peak value of $\mathrm{d}T_{\mathrm{p,id}}/\mathrm{d}R$ at corotation increases as $h_{\mathrm{p}}$ decreases, which is caused by the decreasing softening length (linearly scaling with $h_{\mathrm{p}}$) of the planetary potential that contributes to the density perturbation at the corotation. For that reason the strength of the indirect corotation contribution $T_{\mathrm{c,id}}$ should depend on $h_{\mathrm{p}}$, with the scaling $T_{\mathrm{c,id}}\propto h_{\mathrm{p}}^{-1}$ (Goldreich and Tremaine, 1979), which is indeed the case (see Section 5.1).
For $p=1.5$ (Figure 12b) the peak at the corotation is absent, in agreement with the theoretical expectation (Goldreich and Tremaine, 1979, 1980) for the lack of corotation torque in a $p=1.5$ disc, which has no vortensity gradient. Thus, in this case we assign the full indirect torque due to the disc to $T_{\mathrm{L,id}}$.
Figure 12 also clearly shows that in the framework of our linear calculation the inner disc provides essentially no contribution to $T_{\mathrm{p,id}}$, consistent with analytical expectations. Indeed, $T_{\mathrm{p,id}}$ is determined by the $m=1$ Fourier harmonic $\delta\Sigma_{1}=\Sigma_{1}$ of the surface density perturbation (see equation (36)), which is extremely small in the inner disc for $M_{\mathrm{p}}\ll M_{\mathrm{th}}$ according to Figure 2b. We note in this regard that Crida et al. (2025) did find a considerable non-zero (albeit rapidly oscillating) contribution to $T_{\mathrm{p,id}}$ from the inner disc. This may be somehow related to the higher $M_{\mathrm{p}}=0.08M_{\mathrm{th}}$ that they considered; our $M_{\mathrm{p}}=0.3M_{\mathrm{th}}$ calculation clearly shows small but non-zero $\Sigma_{1}$ in the inner disc (see Figure 3b), likely caused by the growth of disc eccentricity in the inner disc, see Section 3.2.
Both panels of Figure 12 make it clear that excitation of the Lindblad torque $T_{\mathrm{L,id}}$ is not concentrated in the immediate vicinity of $R_{\mathrm{L}}$, as expected from the classical analysis of (Goldreich and Tremaine, 1979, 1980) and mentioned in Section 5.1. In reality, near and outside the $m=1$ OLR, $\mathrm{d}T_{\mathrm{p,id}}/\mathrm{d}R$ exhibits slowly decaying oscillations and $T_{\mathrm{L,id}}$ converges to its final value rather far [^11] The non-locality of tidal coupling at Lindblad resonances is also the reason behind the negative torque density phenomenon studied in Dong et al. (2011a) and Rafikov and Petrovich (2012). from the planet. The origin of these oscillations is analogous to that of the binary torque oscillations in circumbinary discs (Cimerman and Rafikov, 2024) and can again be traced to the global structure of the $m=1$ perturbation in the disc (see Figure 1).
@@ C.2 Direct $m=1$ Lindblad torque $T_{\mathrm{L,d}}$
To compute the direct $m=1$ Lindblad torque component $T_{\mathrm{L,d}}$ we will use the fact that the direct planetary torque on the disc is equal to the direct disc torque on the planet (e.g. see equation (30)). This allows us to use the torque calculation machinery developed by Goldreich and Tremaine (1979), see equations (A7)-(A10) in the Appendix of that work. However, we need to generalize their results, since Goldreich and Tremaine (1979, using their notation) computed the torque exerted by the potential perturbation component $\varphi_{1}$ on the surface density perturbation $\sigma_{1}$ generated by the same potential perturbation $\varphi_{1}$; note that here the subscripts in $\sigma_{1}$ and $\varphi_{1}$ refer to first-order perturbations, not $m=1$ Fourier components. In our case, $\sigma_{1}$ is due to the full perturbing potential — i.e. the sum of the direct and indirect planetary potentials – but the calculation of $T_{\mathrm{L,d}}$ must account only for the torque due to the direct planetary potential acting on this $\sigma_{1}$.
Let us consider a general case when the $\Sigma$ perturbation is caused by a perturbing potential (its $m$-th azimuthal harmonic) $\Phi_{\Sigma}$, while the force enacting the torque is the gradient of some other potential $\Phi_{\mathrm{T}}$. This means that in the equation (A7) of Goldreich and Tremaine (1979) we associate $\varphi_{1}$ with $\Phi_{\mathrm{T}}$, while $\sigma_{1}$ is induced by $\Phi_{\Sigma}$. Propagating this through the derivation in the equations (A7)-(A10) of Goldreich and Tremaine (1979), we find the corresponding Lindblad torque at $m$-th OLR to be
@@
In our case the perturbation of $\Sigma$ is caused by the full planetary potential, including its indirect part, $\Phi_{\Sigma}=-(GM_{\mathrm{p}}/R_{\mathrm{p}})[b^{(1)}_{1/2}(\alpha)-\alpha]$, $\alpha=R/R_{\mathrm{p}}$ for $m=1$ (Goldreich and Tremaine, 1980). To find the direct part $T_{\mathrm{L,d}}$ of the $m=1$ Lindblad torque, we take $\Phi_{\mathrm{T}}$ to be given by only the direct part of this expression, $\Phi_{\mathrm{T}}=-(GM_{\mathrm{p}}/R_{\mathrm{p}})b^{(1)}_{1/2}(\alpha)$. Plugging this into (48) and setting $m=1$, $p=1.5$ (to express $\Sigma(R_{1})$ through $\Sigma_{\mathrm{p}}$ and enable comparison in Figure 13) and $R_{\mathrm{L}}=R_{1}=2^{2/3}R_{\mathrm{p}}$, we obtain
@@
which is independent of $h_{\mathrm{p}}$, in lieu of the definition (15). This is in agreement with the classical theory (Goldreich and Tremaine, 1979, 1980) predicting that the torque excited at an individual Lindblad resonance should be independent of $c_{\mathrm{s}}$ and $h_{\mathrm{p}}$.
@@ C.3 Indirect Lindblad torque $T_{\mathrm{L,id}}$
To find $T_{\mathrm{L,id}}$ we cannot use the symmetry property of the direct torques that we exploited for computing $T_{\mathrm{L,d}}$, see the discussion around equations (31) and (32). Instead, we note that our equation (36) looks identical to the initial torque expression (A7) of Goldreich and Tremaine (1979) if we set $\varphi_{1}=\Phi_{\mathrm{T}}=(GM_{\mathrm{p}}R_{\mathrm{p}}/R^{2})$; this implies that the subsequent result (48) can still be used to compute $T_{\mathrm{L,id}}$ if we adopt this ‘effective’ form of $\Phi_{\mathrm{T}}$ (Fairbairn and Rafikov, 2025), keeping $\Phi_{\Sigma}$ as above. Performing this procedure for a $p=1.5$ disc, we arrive at
@@ C.4 Lindblad torque estimates: verification
We verify our analytical calculation of $T_{\mathrm{L,d}}$ and $T_{\mathrm{L,id}}$ in Figure 13 using our numerical linear calculations for $p=1.5$ (when $T_{\mathrm{c,id}}=0$) and $q=0$, finding good agreement with the analytical results given by equations (50) and (51). The variation of these torque components with $h_{\mathrm{p}}$ is rather minor, as expected. The slight differences for lower $h_{\mathrm{p}}$ are likely caused by the non-locality of disc-planet coupling at the $m=1$ OLR, as noted earlier. Our present results, as well as those of Korycansky and Pollack (1993), Rafikov and Petrovich (2012) and Fairbairn and Rafikov (2025), suggest that the simplifying assumption of the locality of Lindblad torque excitation near the OLR, which was adopted in Goldreich and Tremaine (1979, 1980), may result in some minor inaccuracies.